### v2

# AI-Informed Model Analogs for Subseasonal-to-Seasonal Prediction

Abstract
Subseasonal-to-seasonal (S2S) forecasting is crucial for public health, disaster preparedness, and agriculture, yet it remains particularly challenging to forecast and to understand the modes of predictability on this timescale. We adapt an interpretable AI-informed analog forecasting approach, previously used for longer timescales, to improve S2S analog prediction and understanding of its climate drivers. Using an artificial neural network, we learn a mask of weights to optimize analog selection and showcase its versatility across two prediction tasks: 1) regional continuous prediction of Month 1 midwestern U.S. summer temperatures and 2) classification of Month 1-2 North Atlantic wintertime upper atmospheric winds. The AI-informed analogs outperform traditional analog forecasting approaches, as well as climatology and persistence baselines, for deterministic and probabilistic skill metrics on both climate model and reanalysis data; moreover, this skill gap grows for extreme predictions. Moreover, by using an interpretable-AI framework, we analyze the learned masks of weights to understand the underlying physical processes and find skin temperature and the Northern Hemisphere to be important predictors of North Atlantic wintertime upper atmospheric winds.
: Machine Learning: Earth
1 Introduction
Forecasting on S2S timescales, typically defined as 2 weeks to $\sim$2 months, is vital for public health, disaster preparedness, agriculture, and energy/water management [White:2017aa]. Despite the clear benefits of skillful predictions on these timescales, S2S forecasting remains especially difficult. Often referred to as a ‘predictability desert’ [Robertson:2018ab, Chen:2024aa], S2S forecasts cannot solely rely on the initial atmospheric conditions, as is often done in short-term numerical weather prediction, or on the slow-varying boundary conditions that underpin climate outlooks [Robertson:2018ab, Vitart:2018aa]. Instead, forecasters must integrate information from initial conditions, boundary conditions, and S2S modes of variability, like the Madden Julian Oscillation (MJO) [Zhang:2013aa], to produce skillful predictions [Vitart:2018aa]. Still, on S2S timescales, the strength of these sources of predictability and their teleconnections remain unclear [Merryfield:2020aa, Vitart:2018aa] and skill, e.g. accuracy of summertime surface temperature prediction in North America, remains relatively low [Breeden:2022aa, Pegion:2019aa].
A variety of tools have been used to approach the S2S forecasting challenge. Dynamical models have slowly but steadily improved S2S forecast skill [Peng:2023aa] and data-driven approaches, like fully-AI models, can now forecast phenomena such as the North Atlantic Oscillation (NAO) and MJO at S2S lead times [ling2024fengwuw2sdeeplearningmodel, Chen:2024aa] with similar skill to dynamical models. To further improve forecasts, there has recently been a renewed focus on pinpointing climate states that represent times of enhanced predictability [[, e.g.,]]Mariotti:2020aa,Mayer:2021aa, Albers:2019aa. Identifying these ‘windows of opportunity’ is a potential approach to improve skill on S2S timescales by allowing forecasters to know when forecast uncertainty is high or when they can leverage these times of enhanced predictability for more accurate forecasts [Mariotti:2020aa].
Here, we tackle S2S prediction by combining a variety of these methodologies and employing an AI-informed model analog forecasting approach. Analog forecasting rests on the premise that climate states with similar initial conditions tend to evolve in a consistent manner [[, e.g.,]]Lorenz:1969aa,Zhao:2016aa. By identifying past states resembling current conditions, their subsequent evolution can offer plausible trajectories for future conditions. For a variety of forecasts, from the tropics to the northern high latitudes, analog forecasting has been shown to rival the skill of global climate models [Lou:2023aa, Ding:2019aa, Walsh:2021aa] all while offering several key advantages. Unlike fully-AI models, analog forecasting is intuitive, interpretable, and can uphold physical laws [Rader:2023aa, Ding:2018aa]; moreover, compared to global dynamic climate models, analog forecasting is highly computationally efficient [Ding:2019aa].
Analogs offer an interpretable, physical model that is helpful for diagnosing errors and probing physical drivers, while their fast computational speed allows for the quick generation of ensembles of forecasts. Creating proficient ensembles is a key way to improve skill on S2S timescales [[, e.g.,]]Han:2023aa,Palmer:2004aa,Krishnamurti:1999aa, provide probabilistic forecasts [[, e.g.,]]Mullan:2006aa,Leutbecher:2008aa,Weisheimer:2014aa, and even help explore windows of opportunity [[, e.g.,]]Leutbecher:2008aa,Weisheimer:2014aa—essential on S2S timescales. For instance, with a calibrated ensemble of forecasts, one can use ensemble member agreement as a sign of a lower forecast uncertainty to identify windows of opportunity [[, e.g.,]]Ferranti:2018aa. However, despite these advantages in computation and interpretability, successful analog forecasting hinges on having both a robust library of analogs and a reliable method to identify sufficiently similar past states.
To address this need for a large analog library, we turn to climate models, which have orders of magnitude more climate realizations than we have observational data [Ding:2018aa, McDermott:2016aa]. Yet, even with climate models, finding perfect analogs is impractical—estimates suggest over $10^{30}$ years of data would be needed to match two atmospheric flow stream patterns in just the Northern Hemisphere within observational error [VAN-DEN-DOOL:1994aa]. Hence, determining the conditions that make a climate state an adequately close analog, rather than a perfect one, is crucial. For example, [Ding:2018aa] use regional matching to identify close analogs for seasonal tropical Indo-Pacific Ocean prediction; [esd-13-1437-2022] use global matching for multi-decadal global predictions; and [Wu:2023aa] use area-specific matching for annual-to-multi-year Pacific Decadal Oscillation prediction. These methods for selecting analogs have been shown to work for certain problems, although they demand either a huge library of analogs (as in global matching) or depend on prior knowledge of physical drivers and teleconnections (as in regional or area-specific matching).
Here, we explore an alternative, AI-based spatial weighting approach originally introduced by [Rader:2023aa]. We train a neural network to output a mask of weights that highlights where it is most important for initial conditions to match, such that two states will evolve similarly. Using a learned set of weights to find optimal analogs reduces reliance on prior knowledge and enables investigation of which regions and variables are most essential for two climate states to follow similar future trajectories. This method of optimized analog forecasting was first successfully applied to annual-to-decadal sea surface temperature prediction [Rader:2023aa] and has since been extended to multi-year-to-decadal 2-meter temperatures [fernandez2025multiyeartodecadaltemperaturepredictionusing] and seasonal-annual El Niño-Southern Oscillation (ENSO) predictions [toride2024usingdeeplearningidentify].
Here, we show that this AI-based analog forecasting approach can achieve skill beyond traditional analog methods on S2S timescales while maintaining interpretability and computational efficiency. We highlight the benefits of using AI-based analogs across two prediction tasks: 1) regional continuous prediction of Month 1 midwestern U.S. summer temperatures and 2) classification of Month 1-2 North Atlantic wintertime upper atmospheric winds. Through these predictions tasks we show the AI-based analog approach outperforms traditional analog forecasting approaches, climatology, and persistence on reanalysis data on S2S timescales, exhibiting especially strong performance for extreme temperature prediction. Moreover, via this interpretable AI-forecasting framework we analyze the learned masks of weights to better understand the S2S sources of predictability that underpin this improvement.
2 Methods
2.1 Prediction Tasks
We demonstrate the skill of our analog forecasting approach for the prediction tasks described in Table 1, opting for a varied pair of examples to test the generalizability of the method across different S2S prediction problems. We apply the AI-informed analog approach to both classification and continuous prediction tasks, to different regions, seasons, variables, and lead times. These prediction tasks each have a unique learned mask of weights to optimize the choice of analogs. While we focus on monthly prediction, we also find skill over traditional analog methods for week 3-4 prediction in Task #3, included in the Supplemental Section S1.
Prediction Task 1 2 3 (Supplemental)
Region Midwestern U.S. North Atlantic Southern California
Data Frequency Monthly Monthly Smoothed Daily
Prediction Time Month 1 Month 1-2 Week 3-4
Prediction Season July - Sept. Dec. - Feb. June Week 3 - Sept. Week 3
Input $\to$ Target Skin Temp $\to$ Skin Temp Skin Temp + U250 $\to$ U250 Skin Temp $\to$ Skin Temp
Target Type Single Value Field Single Value
Classification/Value Continuous Tercile Classification Tercile Classification
Table 1: The three prediction tasks.
2.2 AI-Informed Analog Approach
Most traditional analog forecasting methods follow a similar approach: to predict how a certain climate state (referred to as a state of interest or SOI) will evolve, one finds the closest $k$ matches (for $k\geq 1$, in the analog) library. “Closeness” is often measured by minimizing a distance measure, like mean-squared error (MSE), between the SOI and potential analogs either across the entire globe, or across a region of interest. One can then use the trajectories of these closest matches as a prediction for how the SOI will evolve into the future. Rather than predicting the evolution of the whole globe, one often evaluates the analog skill in a specific region of interest, which we refer to as the target region.
We take a similar approach, except here we utilize a soft mask of weights to measure closeness between the SOI and potential analogs. Prior to computing the distance measure between the SOI and each potential analog, we multiply the entire library (after seasonal subsetting) and the SOI by a learned mask of weights (Steps 1 and 2 in Figure 1). This mask, therefore, highlights or dampens the importance of conditions matching in certain areas of the globe for a potential analog to be considered close to the SOI. We then use MSE to compute the closest analogs after weighting, selecting the $k$ closest analogs (Step 3 in Figure 1). Lastly, we use the $k$ closest analogs’ mean evolution (for continuous prediction problems) or majority vote (for classification) in the target region as our final prediction (Steps 4 and 5 in Figure 1).
Figure 1: Schematic of the steps in the AI-informed analog approach: 1) Assemble the SOI and the library of $d$ potential analogs. 2) Multiply all potential analogs and the SOI by the learned mask of weights. 3) Compute the MSE between the weighted SOI and the $d$ weighted potential analogs, and select the $k$ closest analogs (in this example $k=2$). 4) Find the values of the analogs in the target region after the desired lead time. 5) Use the target field of the $k$ closest analogs’ evolution in the target region as the prediction for the SOI (this example is a continuous prediction problem, so the mean value is taken).
2.3 Data
We predict both U250 and skin temperature, relying on output from the Community Earth System Model 2–Large Ensemble (CESM2-LE) [Danabasoglu:2020aa] in order to have a sufficiently large analog library and to learn the weighted mask. As we will show, the AI-informed analog approach produces skillful predictions when evaluated on both CESM2-LE data, in a perfect-model framework, and on ECMWF Reanalysis v5 (ERA5) data [Hersbach:2020aa]. We make use of monthly-mean fields that is resolved at .25° x .24° (natively for CESM2-LE data, and bilinearly interpolated for ERA5 data). For all data sources, we convert the data to anomalies about the climatological seasonal cycle and then to standard deviations across the subsetted season at each grid point. However, between data sources, we handle the anthropogenic effects of climate change slightly differently, as will be discussed next.
2.3.1 CESM2-LE Data
We use monthly CESM2-LE data from 1850-2100 that employs CMIP6 historical and SSP3-7.0 future radiative forcing scenarios [Simpson:2023aa]. We take all 100 members to calculate the ensemble mean, which we subtract from each individual member to both remove the effects of anthropogenic climate change and to convert the data to anomalies from the seasonal cycle. To increase speed and reduce memory load, we then use only a third of the members for training and the analog library. These members are divided between the analog library and SOIs, with fields from 19 members composing the library and fields from 14 members serving as the SOIs (see Table S2 for member details). We partition the SOIs with a 10/2/2 member split for training, validation, and testing respectively.
2.3.2 ERA5 Data
We use ERA5 data from January 1940 to July 2024. We fit and subtract a third-order polynomial at each grid point and each calendar month to define detrended anomalies from the seasonal cycle. The ERA5 data acts as a second test set to evaluate skill on observations.
2.4 Artificial Neural Network to Learn Mask of Weights
The goal of the artificial neural network is to optimize the premise of analog forecasting; namely, that smaller differences in initial conditions lead to smaller differences in future conditions. To optimize this relationship between initial and future conditions, we task the network with predicting how similarly two states will evolve in the target region given their initial conditions. We employ a network that is similar to that of [Rader:2023aa], with minor modifications to its final layers. During each forward pass, the network, depicted in Figure 2, takes two maps as input and uses these initial conditions to predict the similarity of their future states. The SOI map is from the training set and the analog map is randomly selected from the analog library. These maps are both multiplied by a grid of learnable weights (i.e., the mask) of the same size as the inputs, resulting in two weighted maps. The mask is restricted to have a mean of 1 across all weights, such that during training weight is moved between different areas of the globe, but conserved. The MSE between these two weighted maps is calculated, representing the effective similarity of the two initial states for this prediction task, and is passed through a single linear scaling layer. The output of this layer represents the network’s prediction of the MSE between the two maps in the target region after they have evolved (i.e., after the desired lead time). Loss is computed as the MSE between the predicted difference of the targets and the true difference of the targets. Hence, through this weighting process, the network learns a mask that aligns the MSE between two states’ inputs to their MSE after evolution in the target region. This process is repeated for each SOI in the training set. Details of the network setup and hyperparameters can be found in Table S1.
Our network deviates from that of [Rader:2023aa], in that we use a single linear layer instead of multiple dense layers at the end of the network. We restrict the linear layer’s weight to be $\geq 0$ to ensure a monotonically increasing relationship between the MSE of the maps’ weighted inputs and the predicted MSE of the maps’ target regions. This better matches our process for selecting analogs as described in Section 2.2, as we expect that two maps with a smaller weighted MSE will also evolve to have a smaller MSE between their targets. This switch to a single linear layer resulted in a negligible change in skill, but increased network parsimony and training speed.
Figure 2: Schematic of the neural network setup to learn the weighted mask. One SOI and one analog are multiplied by a layer of learnable weights. The MSE between the two weighted inputs is computed and passed through a linear scaling layer. This output represents the predicted difference in the two maps’ targets. Loss is computed as the MSE between the predicted difference of the targets and the true difference of the targets.
2.5 Metrics
We employ deterministic and probabilistic error metrics for each type of prediction task (i.e., classification and continuous prediction). For continuous prediction (Task #1) we compute mean absolute error (MAE) and continuous ranked probability score (CRPS).
MAE is defined as
$MAE=\frac{1}{N}\sum_{i=1}^{N}|f_{i}-o_{i}|$ (1)
where $N$ is the number of samples, $f_{i}$ is the predicted value for sample $i$, and $o_{i}$ is the true value for sample $i$.
CRPS is defined as
$CRPS(F,x)=\int_{-\infty}^{\infty}\left(F(y)-H(y-x)\right)^{2}dy$ (2)
where $F(y)$ is the cumulative distribution function of the forecast, $x$ is the true value, and H(yx)H(y-x) is the Heaviside step function, which is 0 for $y<x$ and 1 for $y\geq x$. CRPS ranges from 0 (for a perfect forecast) to $\infty$.
For classification (Tasks #1 and #3), we compute misclassification rate and the multi-class Brier Score (BS).
Misclassification rate is defined as
$\mathrm{Error\ Rate}=\frac{\mathrm{Number\ of\ Incorrect\ Classifications}}{\mathrm{Total\ Number\ of\ Predictions}}$ (3)
BS is defined as
$BS=\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}\left(f_{ik}-o_{ik}\right)^{2}$ (4)
where $N$ is the number of samples, $K$ is the number of classes, $f_{ik}$ is the predicted probability for class $k$ for sample $i$, and $o_{ik}$ is the true value (1 if the true class is $k$, otherwise 0). BS ranges from 0 (for a perfect forecast) to 2.
We convert all types of error to skill scores by comparing them to the error of a climatological forecast:
$\mathrm{Skill\ Score}=1-\frac{Error}{Error_{climatology}}$ (5)
All skill scores are strictly $\leq 1$, with a skill score of 1 indicating perfect skill and a skill score of 0 indicating equal skill to a climatological forecast. A negative skill score indicates worse skill than a climatological forecast.
2.5.1 Baselines
We include a regional and a global analog baseline in addition to persistence, climatological, and random baselines to evaluate the relative skill of the learned mask approach. To create a global baseline, we select analogs by matching conditions over the entire globe (equivalent to a weighted mask of 1s everywhere). We create a regional baseline by selecting analogs via matching conditions only in the target region (equivalent to a weighted mask of 1s in the target region and 0s everywhere else). The 90th percentile random baseline is formed by repeating the prediction for all SOIs using random analogs 100 times and selecting the 90th percentile of best predictions.
3 Results
3.1 Month 1 Temperature Extremes Over the Midwestern U.S.
[…]
Figure 3: The learned mask for Task #1, midwestern U.S. summer temperatures. The cyan box outlines the target region.
The learned mask modestly outperforms all baselines (Figure 4), with MAE skill increases of 17% and 51% and CRPS skill increases of 5% and 48% tested on CESM2-LE and ERA5, respectively. All skill scores peak at 50 analogs for both CESM2-LE and ERA5 data, except for CESM2-LE CRPS, which peaks at 100 analogs. This overall improvement in temperature forecasting on S2S timescales is important, however, better prediction of extreme temperatures in particular has an outsized impact on enhancing agricultural production, public health, and energy management [Domeisen:2022aa]. Thus, we focus on assessing the AI-based analog’s ability to predict extreme temperatures. Here, we utilize a discard plot, in which we progressively discard samples with lower extremity to visualize how MAE skill changes for more extreme samples. We denote extremity simply as the absolute value of the prediction, i.e. a measure of how far from climatology the prediction is.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Skill scores for a) CESM2-LE MAE, b) CESM2-LE CRPS, c) ERA5 MAE, and d) ERA5 CRPS for Month 1 midwestern U.S. temperatures.
In Figure 5, we show the discard plot with ERA5 data using an ensemble of 50 analogs. The AI-based analogs exhibits a marked increase in skill for samples with more extreme predictions. As our skill score is defined relative to climatology, it may be unsurprising that the AI-informed analogs would have lower relative error on more extreme events. However, this is not the case for the regional baseline, where there is only a slight increase in skill for the most extreme samples. This analysis highlights how the skill gap between AI-informed analogs and traditionally-selected analogs widens for more extreme temperature events. Moreover, as we use predicted extremity to discard samples, this information is available a priori, allowing forecasters to better understand when the analog ensemble is likely to perform best and building trust in its more extreme predictions. This behavior also holds for CESM2-LE data (Figure S6).
Figure 5: Discard plot based on predicted extremity with an ensemble of 50 analogs for midwestern U.S. summer temperatures, testing on ERA5 data. Data with the lowest extremity is progressively discarded, with the x-axis showing the percentage of data remaining.
3.2 Month 1-2 Winds in the North Atlantic
Next, we explore the learned mask’s ability to perform grid-point classification of upper atmospheric winds in the North Atlantic ($25^{\circ}-48^{\circ}\mathrm{N},\ 0^{\circ}-80^{\circ}\mathrm{W}$) and probe the mask itself to better understand the relative importance of different areas for successful prediction. At each grid point in the target region (rather than averaging across the target region), we classify the 250 hPa zonal wind (U250). The three target classes are formed by splitting the target temperatures into terciles, ensuring all classes are equally sized. Terciles for classifying the analog library are determined using the data within the analog library and terciles for the test set are defined based on the data in the test set to limit the impact of CESM2-LE biases relative to ERA5. We make predictions for December-January and January-February, using the learned mask in Figure 6. We chose to examine the winter winds, as the jet stream variability in this region is largest during this time [[, e.g.,]]Hall2017. In this case, we select analogs using both U250 and surface temperature as inputs. Therefore, we learn a unique mask for each field, although, importantly, these masks are learned together by the network. As with Tasks #1 and #3, we include the predictand as an input variable both for predictive power (e.g., to capture jet stream information) and to indirectly allow for the encoding of persistence. Here we also include skin temperature as an additional input variable, as it encodes signal from ENSO and the MJO which can drive North Atlantic atmospheric variability [[, e.g.,]]ENSO_NAO,NAO,MJO_NAO_teleconnections, TheConsistencyofMJOTeleconnectionPatternsAnExplanationUsingLinearRossbyWaveTheory. Thus, despite the drawback in increased memory load, we find including both U250 and skin temperature are useful inputs, as we discuss further in 3.2.1.
Figure 6: The learned mask for Task #2, North Atlantic winter U250 classification. The cyan box outlines the target region. The dark green boxes outline areas of high weight in the mask.
We evaluate skill at each grid point in the whole field (e.g., Figure S7), summarizing these with mean skill scores over the target region (Figure 7). The learned mask outperforms all baselines for this grid-point-by-grid-point classification, with accuracy skill increases of 6% and 1% and BS skill increases of 19% and 13% tested on CESM2-LE and ERA5, respectively. Skill peaks at 400 analogs for both CESM2-LE and ERA5 data, except for CESM2-LE classification, which peaks at 800 analogs. The numbers of analogs in the ensembles are much higher than in Task #1, since we have moved from a continuous prediction problem to a 3-class classification problem. With a continuous prediction problem, if the number of analogs in the ensemble is too high, the ensemble mean will converge to climatology. An example of this regression to the mean can be found in Figure S5. With a majority vote classification problem, however, the ensemble can be larger without this issue, as the majority vote will not converge to climatology.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Skill scores for a) CESM2-LE accuracy, b) CESM2-LE BS, c) ERA5 accuracy, and d) ERA5 BS for Month 1-2 North Atlantic U250 classification.
3.2.1 Physical Interpretation of the Learned Mask
Using this interpretable AI-informed analog approach, we can analyze the learned mask to better understand the physical drivers behind analog predictions. Here we discuss three areas of high weight in the mask, shown in Figure 6: 1) SSTs in the eastern tropical Pacific, 2) SSTs in the Phillipine Sea, and 3) U250 in the Northern Hemisphere. Area 1 resembles the canonical ENSO region. We find this reflected in the analog selection, with selected analogs generally have more similar ENSO states than a random selection (Figure S10), as measured by Niño-3.4 indices [AVerifiedEstimationoftheElNioIndexNio34since1877]. This is in line with work that finds ENSO impacts the North Atlantic region through its influence on the North Atlantic Oscillation (NAO)—the most prominent pattern of atmospheric variability in the region [ENSO_NAO, NAO]. Area 2 resembles a prominent Rossby wave source region, which is a known source of MJO teleconnections to the North Atlantic [MJO_NAO_teleconnections, TheConsistencyofMJOTeleconnectionPatternsAnExplanationUsingLinearRossbyWaveTheory]. Lastly, Area 3 highlights the Pacific jet exit region of the subtropical jet stream. The region of high weight over the Eastern Pacific and Mexico aligns with the mean position of the subtropical jet, and both its strength and the two North–South shifted branches, shown in smaller green boxes in Figure 6, have been found to vary with the NAO state [JET_NAO].
We also directly probe the learned mask to better understand the relative importance of initial conditions in different areas for successful analog prediction. We do so by ablating the mask, i.e. setting the weights to 0, and observing changes in skill. We analyze changes in BS skill with CESM2-LE data and a 400 analog ensemble. We test on CESM2-LE data rather than ERA5 because the impacts on BS are small, and thus, there are too few samples to draw meaningful conclusions from ERA5 data. We employ three ablation methods: 1) threshold ablation, where we increase mask sparsity by setting weights to 0 if they are below either the 40th, 80th, or 90th percentile or incentivizing sparsity during training itself by adding constrained inverse $L_{2}$ regularization (see S2.2 for details), 2) Ablating entire fields (e.g. temperature or U250), and 3) ablating specific regions (e.g. the Northern Hemisphere). Masks for examples of these ablation methods are shown in Figure 8. All of these ablation methods, except constrained inverse $L_{2}$ regularization, are performed after the mask has already been learned.
Figure 8: Examples of masks with each of the ablation methods: (a) Threshold ablation, (b) Ablating entire fields, and (c) Ablating specific regions.
We focus on a 400 analog ensemble, as this is the number of analogs for which skill peaks (Figure 9), although the general trends remain similar across ensemble size (Figure S11). We find a slight improvement in skill when we increase mask sparsity by thresholding or by introducing constrained inverse $L_{2}$ regularization. This increase in skill with a sparser map is consistent with [Rader:2023aa], who found a slight improvement in skill for multi-year predictions using a $\sim 95\%$ percentile threshold. However, when we test increasing the sparsity for shorter timescales (e.g. Task #3), we find minimal change in observation skill and a slight decrease in probabilistic model skill (Figure S9). Considering field ablation, temperature appears to be the more important of the two fields, although ablating either the temperature field or the U250 field results in a significant decrease in skill, highlighting the importance of both for identifying skillful analogs. While all ablation methods besides increasing sparsity decreases skill, ablating the Northern Hemisphere, both fields, and ocean temperatures result in the largest drop in skill.
Figure 9: BS skill for different ablation methods evaluated on CESM2-LE data.
4 Conclusions
We demonstrate how an AI-informed model analog forecasting approach, previously only shown to be skillful on seasonal-to-decadal timescales, can also produce skillful deterministic and probabilistic subseasonal-to-seasonal predictions. We showcase this approach’s improvement over climatological, persistence, and traditional analog forecasting methods in both a perfect model framework and with reanalysis data for classification and continuous prediction tasks. For example, we find 49% and 19% increase in probabilistic skill testing on observational data for Tasks #1 and #3, respectively. While skill relative to climatology remains modest, especially for shorter timescales (e.g., Task #1 and Task #3), this is typical for S2S prediction. For example, the Seasonal-to-Multiyear Large Ensemble (SMYLE) prediction system, which similarly utilizes CESM2, finds zero or minimal skill (ACC <.3<.3) when predicting month 1-3 summer Midwestern U.S. temperatures [smyle].
Moreover, the interpretability of this AI-based approach enables users to explore the learned mask to gain insight into the relative importance of different areas of the globe for successful prediction. We perform an analysis of the learned mask for North Atlantic U250 classification, assessing which variable fields and areas of the globe are comparatively more important for a successful analog forecast. This type of analysis can help identify key initial conditions that most influence climate state evolution on S2S timescales, guiding both future model development and observational prioritization.
The AI-informed analog ensembles additionally provides improved prediction of extremes compared to traditional analog forecasting methods. Since extreme temperature events have a disproportionate impact on human health, agriculture, and energy/water management, improving their prediction is essential for mitigating the most dire consequences [annurev:/content/journals/10.1146/annurev-earth-071719-055228].
Like previous work [[, e.g.,]]Rader:2023aa,toride2024usingdeeplearningidentify,fernandez2025multiyeartodecadaltemperaturepredictionusing, we leverage a model analog approach rather than rely on observations alone to form our analog library. While we use a single model (CESM2) to form our analog library, [fernandez2025multiyeartodecadaltemperaturepredictionusing] experiment with using multiple models on multi-year-to-decadal timescales to both form the analog library and to train their neural network. Including more climate models or, specifically for S2S timescales, extended-range AI or dynamical weather forecasts [[, e.g.,]]lang2024aifsecmwfsdatadriven,Vitart:2022 may similarly boost skill on these shorter timescales.
Further, although we have shown that this method produces skillful S2S predictions on near observational data, the approach is still limited by the model’s, in this case CESM2’s, biases [[, e.g.,]]Pang:2024aa,Wei:2021aa,Woelfle:2019aa. Future work could also explore how to best incorporate near-observational data into the AI-based analog forecasting approach. For example, [fernandez2025multiyeartodecadaltemperaturepredictionusing] found that transfer learning from models to reanalysis improved analog prediction skill; a similar approach could be useful for learning masks on S2S timescales as well. Moreover, reanalysis data could also be incorporated into the analog library itself, perhaps offering even more realistic climate trajectories than those from a library composed solely of model data. These additions to the analog library and/or training set could further advance AI-based analog S2S forecasts, especially for regions and seasons where model biases are more pronounced.
5 Acknowledgments
J.B.L. acknowledges support from NOAA grants #NA22OAR4310621 and #NA19OAR4590151. The authors wish to thank Martin Fernandez for helpful insights for this research.
6 Data and Code Availability
CESM2-LE data is available at https://www.earthsystemgrid.org/dataset/ucar.cgd.cesm2le.output.html [Danabasoglu:2020aa]. ERA5 data is available at https://cds.climate.copernicus.eu/datasets [Hersbach:2020aa]. Code used to perform the analysis and generate the figures is available at https://github.com/jlandsbe/S2S.git.
[]Supplemental Information for AI-Informed Model Analogs for S2S Prediction
S1 Week 3-4 Windows Southern California Predictions
S1.1 Daily Data
We employ a 7-day sliding window to smooth daily CESM2-LE and ERA5 data, using a backward moving average for input data and a forward moving average for target data. All smoothed daily data is regridded via bilinear interpolation to 2.5° x 2.5° resolution. We use a coarser resolution for the daily data to reduce the memory load, as our analog library of daily data climate maps is $\sim 10\times$ larger than the library of monthly data. This data is similarly converted to anomalies about the seasonal cycle and then to standard deviations at each grid point. For the daily CESM2-LE data, we subtract the linear trend from each calendar day at each grid point. We include a shorter timespan of data from each member (1850-1949) than the monthly data, as each daily-data year contains more than $30\times$ the amount of samples. The analog library is composed of fields from the first 5 members, while fields from the next 4 members (with a 2/1/1 training/validation/testing split) make up the SOIs (see Table S3 for member details). For daily ERA5 data, we use dates between 1942-2023, fitting and subtracting a third-order polynomial at each grid point and each calendar day to define detrended anomalies.
S1.2 Week 3-4 Results
We also assess the short-range S2S skill of the AI-based analog approach by classifying Week 3-4 Southern California ($32^{\circ}-37^{\circ}\mathrm{N},\ 116^{\circ}-121^{\circ}\mathrm{W}$) summer temperatures (Task #1). The three target classes (cold, neutral, and warm) are formed as in Task #2, where target temperatures are split into terciles, ensuring all classes are equally sized. We predict each 2-week period from the third week of June through the third week of September, using the learned weights in Figure S1. This mask exhibits weights that are distributed globally, yet unevenly, with noticeably increased weight around the western U.S. as well as in the North Pacific. The more diffuse weighting pattern, as compared to Tasks #1 and #2, likely reflects the noisier synoptic patterns present in the daily data used to train this mask.
Figure S1: The learned mask for Task #3, Southern California summer temperature classification. The cyan box outlines the target region.
With the learned mask, MAE and BS skill scores exceed other baselines, with increases in MAE skill of 16% and 18% and CRPS skill of 4% and 71% when testing on CESM2-LE and ERA5 data, respectively (Figure S2). With the CESM2-LE test set, the highest skill is reached at 2000 analogs, while for ERA5, the skill score peaks at 1500 analogs.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure S2: Skill scores for a) CESM2-LE accuracy, b) CESM2-LE BS, c) ERA5 accuracy, and d) ERA5 BS for Week 3-4 Southern California temperature classification.
We diagnose whether the analog ensembles can offer insights into windows of opportunity via discard plots. Figure S3 uses ERA5 data and a 1500-analog ensemble to show the change in accuracy skill from a climatological forecast as samples with lower ensemble agreement are discarded (for CESM2 data see Figure S4). Here, ensemble agreement is computed as the fraction of ensemble members that agree on the majority prediction. We see that over all samples the mask offers just over a $4\%$ improvement in accuracy relative to climatology, but this improvement grows essentially monotonically to over $9\%$ for the $\sim 25\%$ of samples with the highest ensemble agreement. This is not the case with a global mask’s ensemble, which does not exhibit as precipitous of an increase in accuracy skill score and actually decreases in accuracy until the $\sim 50\%$ cutoff mark.
Figure S3: Discard plot based on ensemble agreement for Week 3-4 Southern California temperature classification using 1500 analogs, testing on ERA5 data. Data with the lowest ensemble agreement is progressively discarded, with the x-axis showing the percentage of data remaining.
Figure S4: Discard plot for Week 3-4 Southern California temperature classification using 1500 analogs, testing on CESM2-LE data. Data with the lowest ensemble agreement is progressively discarded, with the x-axis showing the percentage of data remaining. Unlike when testing on reanalysis data, global ensemble agreement serves as a reasonable metric for forecast uncertainty. However, the learned-mask ensemble agreement still showcases a greater decrease in error for states with higher ensemble agreement.
S2 Neural Network
S2.1 Hyperparameters
We use the following parameters for the neural network:
Parameter Value
Optimizer Adam
Learning rate .0001
Batch size 32
Loss function Mean squared error
Validation batch size 1000
Early stopping patience 30
Early stopping minimum delta 0.0001
Table S1: Hyperparameters used in the model training.
S2.2 Constrained Inverse $L_{2}$ Regularization
To increase mask sparsity, we implement constrained inverse $L_{2}$ regularization. We do so to compare how post-hoc thresholding and learned sparsity compare in terms of model performance. We add the following term to the loss function:
$\frac{\lambda_{2}}{\sqrt{\sum_{i=1}^{n}w_{i}^{2}}}$ (6)
where $\lambda_{2}$ is the regularization strength, $w_{i}$ is the weight of the $i$th grid point, and $n$ is the total number of grid points in the weighted mask.
This term is restricted during training such that $\sum_{i=1}^{n}w_{i}=n$. With this constraint, the regularization term is maximized (high loss) when $\forall i,\,w_{i}=1$ and minimized when, for some $j$, $w_{j}=n$ and $w_{i\neq j}=0$. Thus, this term promotes having more disparate weight values, with some weights of very high values and some weights of very low values. We generate the mask in Figure S8 for the North Atlantic (Task #3), using this regularization term with $\lambda_{2}=100$.
S3 CESM2-LE Members
For monthly CESM2-LE data, we use the ensemble members listed in Table S2, while for daily CESM2-LE data, we use the ensemble members listed in Table S3. The first four numbers of the member names correspond to the chosen model start dates (varying initial climate conditions), while the last three numbers indicate the realization (small perturbations to initial conditions). While we use a mixture of ocean initializations, these differences have been found to not impact S2S prediction five centuries beyond their initialization date [Arcodia_2023].
Monthly Member Type Members
Analog Library Members 1301.020, 1301.019, 1301.018, 1301.017, 1301.016, 1301.015, 1301.014, 1301.013, 1301.012, 1301.011, 1281.020, 1281.019, 1281.018, 1281.017, 1281.016, 1281.015, 1281.014, 1281.013, 1281.012
SOI Train Members 1301.010, 1301.009, 1301.008, 1301.007, 1301.006, 1301.005, 1301.004, 1301.003, 1301.002, 1301.001
SOI Validation Members 1281.010, 1281.009
SOI Test Members 1281.001, 1281.002
Table S2: List of monthly ensemble members used for the analog library, training SOIs, validation SOIs, and testing SOIs.
Daily Member Type Members
Analog Library Members 1231.012, 1251.013, 1251.014, 1281.015, 1281.016
SOI Train Members 1301.017, 1301.018
SOI Validation Members 1281.017
SOI Test Members 1251.019
Table S3: List of daily ensemble members used for the analog library, training SOIs, validation SOIs, and testing SOIs.
S4 Figures
Figure S5: With regression problems, as the number of analogs approaches the total library size, the prediction becomes more and more similar to a climatological prediction. In this example the library size is $\sim 10000$, where the skill is 0.
Figure S6: A discard plot for an ensemble of 50 analogs for midwestern U.S. summer temperature regression, testing on CESM2-LE data. Data with the lowest extremity is progressively discarded, with the x-axis showing the percentage of data remaining. While the regional mask performs relatively better on extreme predictions for CESM2-LE compared to ERA5 data, there is still a much larger improvement for extreme predictions with the learned mask.
Figure S7: In Task #2 a field of values rather than the average value across the region of interest is predicted. This map shows the average skill across SOIs when using a 400 analog ensemble on ERA5 data. At almost every grid point, the learned mask outperforms the climatological prediction.
Figure S8: Using inverse $L_{2}$ regularization with $\lambda_{2}=100$ for the North Atlantic (Task #2) results in a sparser map.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure S9: Skill scores using a 90th-percentile-thresholded mask for a) CESM2-LE accuracy, b) CESM2-LE BS, c) ERA5 accuracy, and d) ERA5 BS for Week 3-4 Southern California temperature classification. There is little difference in skill between the 90th-percentile-thresholded mask and the learned mask, although there is a slight decrease in Brier Skill Score (BS) for the 90th-percentile-thresholded mask.
Figure S10: Difference in average Niño-3.4 index between the 400 best selected analogs and a random selection of 400. The analogs had a more similar mean Niño-3.4 index compared to the random selection.
Figure S11: Expanded table, showing changes in BS skill with different ablation methods, for 100-800 analogs